Erin's age is 3 times Warren's. In 4 years she will be twice as old as he will be. How old is each now?
Answer by ankor@dixie-net.com(22740)

Let E = Erin's age now
Let W = Warren's age now
:
Write an equation for each statement:
:
"Erin's age is 3 times Warren's."
E = 3W
:
" In 4 years she will be twice as old as he will be."
E + 4 = 2(W + 4)
E + 4 = 2W + 8
E = 2W + 8 - 4
E = 2W + 4
:
How old is each now?
:
Substitute 3W for E in the above equation
3W = 2W + 4
3W - 2W = 4
W = 4 yrs old now
and
E = 3(4)
E = 12 yrs old now
:
:
Check solutions in the statement:
"In 4 years she will be twice as old as he will be."
12 + 4 = 2(4 + 4)